# Robust Radiotherapy Planning with Spatially Based Uncertainty Sets

### v1

@@ 2.2.2 Compactly Reformulating The Robust Problem
In this section, we present an exact reformulation of (2.2) for the case where $\mathcal{U}=\mathcal{U}_{\text{SB}}$ with a polynomial number of constraints with respect to the size of the nominal formulation (2.1).
@@
Lemma 2 (Corazza 1999, Proposition 3.2).
@@
For $S\subseteq\mathbb{R}^{n}$, denote the projection of set $S$ onto index set $I\subseteq[n]$ by $P_{I}(S)$, so that $P_{I}(S)=\left\{v_{I}\;\left|\;\;v\in S\right.\right\}$. For simplicity, for the singleton $I=\{i\}$, the projection will be denoted by $P_{i}(S)$. The following is a straightforward observation that is used in reformulating (2.2).
@@
Problem (2.2) is now specialized and reformulated for the uncertainty set $\mathcal{U}_{\text{SB}}$. Henceforth, it is assumed that $\mathcal{U}_{\text{SB}}\neq\emptyset$. Starting with the the first set of constraints (3c), and using Observation 3, for each $v\in T$, the inner minimization in this constraint is given by
@@
Proof.Part (i): $P_{v}(\mathcal{U}_{\text{SB}})=[\underaccent{\bar}{\phi}_{v},\bar{\phi}_{v}]$ follows from that fact that $\mathcal{U}_{\text{SB}}$ is a convex and compact set and from Observation 3, which together imply that the projection of $\mathcal{U}_{\text{SB}}$ onto $v$ is a closed interval $[\underaccent{\bar}{\phi}_{v},\bar{\phi}_{v}]$.
@@
Two dimensional examples of $\mathcal{U}_{\text{SB}}$ are graphically illustrated in Figure 3. The two-dimensional case will bear a special importance compactly reformulating the robust problem (2.2). Next, we reformulate the second set of constraints (3d). Again, due to Observation 3, for each $v,u\in T$, the inner maximization is
@@
[…]
@@
Proof. Let $u,v\in T$, and for convenience, define also $\tilde{\mathcal{U}}_{u,v}=\{(\phi_{v},\phi_{u}):-\gamma_{vu}\leq\phi_{v}-\phi_{u}\leq\gamma_{uv}\}$, to show that $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})\subseteq\mathcal{U}_{u,v}$. The contractive property of the projection onto a plane, as a convex set, implies that $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})\subseteq P_{u}(\mathcal{U}_{\text{SB}})\times P_{v}(\mathcal{U}_{\text{SB}})$. Then, by Proposition 4-(i) and the definition of $\mathcal{U}_{\text{SB}}$, it follows that
@@
Following the results of Propositions  4-5, formulation (2.2) with $\mathcal{U}=\mathcal{U}_{\text{SB}}$ can now be compactly reformulated with a finite number of constraints. In particular, in this reformulation, no variables are added and the number of constraints of this formulation is $2|T|$ times that of the nominal problem (2.1), thereby remaining polynomial in the size of the nominal formulation. Consider the formulation
$\underset{\begin{subarray}{c}{\bf x}\in\mathbb{R}^{n}_{+},\underaccent{\bar}{d}\in\mathbb{R}\end{subarray}}{\text{maximum}}\hskip-10.0pt$ $\underaccent{\bar}{d}$ (10b)
subject to $\underaccent{\bar}{\phi}_{v}d_{v}(x)\geq\underaccent{\bar}{d},$ $v\in T$ (10c)
$\bar{\phi}_{v}d_{v}(x)-\mu\max\{\bar{\phi}_{v}-\gamma_{vu},\underaccent{\bar}{\phi}_{u}\}d_{u}(x)\leq 0,\hskip-10.0pt$ $(u,v)\in T^{2}:u\neq v$ (10d)
$\min\{\underaccent{\bar}{\phi}_{u}+\gamma_{vu},\bar{\phi}_{v}\}d_{v}(x)-\mu\underaccent{\bar}{\phi}_{u}d_{u}(x)\leq 0,$ $(u,v)\in T^{2}:u\neq v$ (10e)
$\ d_{v}(x)\leq\bar{d}_{k},$ $k\in[K],v\in H_{k}.$ (10f)
The following theorem establishes the correctness of this formulation, and specifically, demonstrates that its optimal solution set coincides with that of (2.2).
@@
[…]
@@
Proof. The reformulation of constraints (3c) is a straightforward consequence of (5) together and Proposition 4. Constraints (3d) are reformulated based on (8) to a maximization over $P_{\{v,u\}}(\mathcal{U}_{\text{SB}})$ in place of a maximization over $\mathcal{U}_{\text{SB}}$. Due to Assumption 1, this set is given by the result of Proposition 5. Further, since the objective is linear, and since, for each $u,v\in T$, $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})$ is a bound polyhedral set, then by the fundamental theorem of linear programming, there is an optimal $(\phi_{u},\phi_{v})$ that is an extreme point of the set $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})$. Moreover, since $\phi_{v}$ and $\phi_{u}$ have, respectively, a positive and a negative coefficient in (3d), this implies that the points in $\text{ext}(P_{\{u,v\}}(\mathcal{U}_{\text{SB}}))$ are dominated by the points in
@@
that is, for each pair $v,u\in T$, constraint (3d) is satisfied if and only if its left-hand side is nonpositive for these two points in $\mathcal{U}^{*}_{u,v}$. ∎
In Section 3, we will address the computational aspect of solving (2.2.2), specifically, how to deal with the potentially very large number of homogeneity constraints (10d) and (10e).
@@ 3 Solving Large-scale RO
The radiation therapy problems that we consider involve tens of thousands of tumor voxels and millions of healthy tissue voxels. Accordingly, formulation (2.2.2) may have hundreds of millions of constraints. Therefore, it may prove difficult to solve or even store this problem on most computers. In order to address this challenge, we propose an algorithm that dynamically generates constraints of (2.2.2) as needed.
@@
1 Phase I: Alternate between iteratively generating at most $n_{H}$ constraints  (10f) to each OAR until a feasibility tolerance $\tau_{\operatorname{Infs},1}$ is achieved and the objective function value does not decrease by more than a factor of $\tau_{z}$, and appending at most $n_{S}$ homogeneity constraints (10d) (10e) to set $S\subseteq T^{2}$. If none of the latter constraints are generated switch to Phase II;
Phase II: The same as Phase I, with feasibility tolerance of $\tau_{\operatorname{Infs},2}<\tau_{\operatorname{Infs},1}$ for constraints  (10f). Exit once no additional constraints can be generated.
@@
A detailed version of the algorithm in pseudocode is given as Algorithm 3 in Appendix A.1. The algorithm is initialized with beamlet configuration $x_{0}$, typically a part of a previously computed solution or an all-ones vector multiplied by some normalized dose. Based on this initial configuration, each OAR is initialized with the most violated $n_{0}$ constraints. This initialization, and the initial emphasis on generating OAR constraints (10f) in the algorithm, ensures that the optimization problem is bounded from the start or that it becomes bounded after a few iterations in phase I of the algorithm. The algorithm consists of two phases. Each one of them alternates between generation of OAR constraints (10f) and (10d) (10e). The first phase involves an OAR constraint infeasibility tolerance $\tau_{\operatorname{Infs},1}$ that is larger than the tolerance applied in the second phase, $\tau_{\operatorname{Infs},2}$. Intuitively, the larger tolerance in the first phase it intended to ensure both that the subproblems become bounded fairly quickly and to delay generation of OAR constraints that may be inactive at optimality until the homogeneity constraints are nearly satisfied.
@@ 4 Dose-Volume Constraints
In general, dose-volume constraints can be formulated by introducing additional auxiliary binary variables into formulation (2.2.2), turning it into the mixed-integer program,
@@
Lemma 7, together with the fact that there exists a value of $\beta>0$ for which $\sum_{v\in H_{K}}y_{v}(\beta)=0$, implies that for every value of $\Theta\geq 0$, it is possible to find a minimal value of $\beta$ such that (12c) is satisfied. Motivating our method is both monotonicity of $\left|\left|y(\beta)\right|\right|_{1}=\sum_{v\in H_{K}}y(\beta)_{v}$, following Lemma 7, and that the one-norm is a convex approximation of the zero-“norm” ($\left|\left|y(\beta)\right|\right|_{0}$) , which is the number of nonzero components of the deviation variable vector $y$. The one-norm is effectively used to promote sparsity in large-scale machine learning and compressed sensing applications (see, for example, Donoho and Elad (2003)). We expect that solving (13) with appropriate $\beta$ will approximate the problem with dose-volume constraints (11f) reasonably well. Our goal is to find the minimal value of $\beta$ for which the optimal solution of (13) satisfies (11f). To achieve this , LP (13) is solved parametrically, starting from a lower bound on $\beta$ and traversing optimal bases for the given subsets of constraints defining the reduced problem. This algorithm is presented as Algorithm 2.
@@ 5 Experiments
In our experiments, we focus on a particular case study based on real brain data of Patient 4 from the TCIA archive (Clark et al. 2013), referred to as the Brain dataset in Table 1. Due to the very large number of constraints in (2.1), and especially in (2.2.2), for this dataset, we also use a smaller liver imaging dataset (Craft et al. 2014) to evaluate the computational performance of our dynamic generation of OAR and homogeneity constraints in Algorithm 3. FMISO-PET images are not available for this smaller dataset, so the radiosensitivity data are created synthetically using the following formula for $v\in T$: $\phi_{v}=0.85+0.15\times\frac{\max_{u\in T}\{f_{u}\}-f_{v}}{\max_{u\in T}\{f_{u}\}-\min_{u\in T}\{f_{u}\}}$, where vector $f$ follows a multivariate normal density over the PTV voxels, with the mean set as the average voxel x-y-z coordinates, and the covariance proportional to the covariance of the x-y-z coordinates (scaled by 0.2). Details of both datasets are given in Table 1.
@@
$\left|T\right|$ $k$ Name $\left|H_{k}\right|$ $\bar{d}_{k}$ bounds [^1] Based on (ICR 2010)
@@ 5.1 Computational Performance
Our implementation uses the Julia programming language, version 1.6.2, the JuMP package (Lubin and Dunning 2015), and the Gurobi solver, version 9.1.2. The experiments are run on 40 core Linux servers each equipped with Intel Xeon E5-2650, 25 MB cache, 2.3 GHz base frequency CPUs, and 128 GB of RAM. Algorithm 1 (appearing in greater detail as Algorithm 3 in Appendix A.1) is invoked for solving the robust model (2.2.2) and the nominal model (2.1), with parameter values $\tau_{\operatorname{Infs},1}=10$, $\tau_{\operatorname{Infs},2}=0$, $\tau_{z}=10^{-2}$, and $n_{H}=400$. The homogeneity parameter is set to $\mu=1.15$ for Liver instances and $\mu=1.1250$ for Brain instances for the nominal experiments, and $\mu=1.1875$, which obtains a similar homogeneity and biologically-adjusted dose in the nominal scenario. Table 2 shows the running times achieved when solving the nominal model (2.1), which involves only $\left|T\right|$ homogeneity constraints. The results show that dynamic constraint generation is not needed to solve the nominal model but with careful tuning of the constraint generation algorithm, it can improve on the running time required to solve the full model (also yielding significant savings in terms of memory consumption), especially for larger data instances such as Brain. Table 3 shows the running times when applying the robust model to both datasets. The parameters of the uncertainty set are set to $\delta=0.04$ and $\gamma=0.04$ for Liver, and $\delta=0.08$ and $\gamma=0.04$ for Brain. Here, it is evident that dynamic constraint generation is necessary in order to be able to solve both models when the number of homogeneity constraints is $O(\left|T\right|^{2})$. However, dynamically generating all violated constraints may result in excessive running times. In the example shown in Table 3, dynamically generating all violated constraints in each iteration results in a running time of nearly four hours for the Brain instance. In preliminary experiments, with a different choice of parameters than experimented with here, when generating all violated constraints, we have observed cases where the running times took more than 48 hours to solve, and cases memory consumption exceeded its limits.
Table 2: Running time in seconds and number of constraints generated in solving the nominal model (2.1). Here, constraint generation, if indicated, implies that all violated constraints are generated in each iteration.
@@
Table 3: Running time in seconds and number of constraints generated in solving (2.2.2) using Algorithm 3.

### v2

@@ 2.2.2 Compactly Reformulating The Robust Problem
In this section, we present an exact reformulation of (3) for the case where $\mathcal{U}=\mathcal{U}_{\text{SB}}$ with a polynomial number of constraints with respect to the size of the nominal formulation (1).
@@
Lemma 2 (Corazza, 1999, Proposition 3.2).
@@
For $S\subseteq\mathbb{R}^{n}$, denote the projection of set $S$ onto index set $I\subseteq[n]$ by $P_{I}(S)$, so that $P_{I}(S)=\left\{v_{I}\;\left|\;\;v\in S\right.\right\}$. For simplicity, for the singleton $I=\{i\}$, the projection will be denoted by $P_{i}(S)$. The following is a straightforward observation that is used in reformulating (3).
@@
Problem (3) is now specialized and reformulated for the uncertainty set $\mathcal{U}_{\text{SB}}$. Henceforth, it is assumed that $\mathcal{U}_{\text{SB}}\neq\emptyset$. Starting with the the first set of constraints (3b), and using Observation 3, for each $v\in T$, the inner minimization in this constraint is given by
@@
Part (i): $P_{v}(\mathcal{U}_{\text{SB}})=[\underaccent{\bar}{\phi}_{v},\bar{\phi}_{v}]$ follows from that fact that $\mathcal{U}_{\text{SB}}$ is a convex and compact set and from Observation 3, which together imply that the projection of $\mathcal{U}_{\text{SB}}$ onto $v$ is a closed interval $[\underaccent{\bar}{\phi}_{v},\bar{\phi}_{v}]$.
@@
Two dimensional examples of $\mathcal{U}_{\text{SB}}$ are graphically illustrated in Figure 3. The two-dimensional case will bear a special importance compactly reformulating the robust problem (3). Next, we reformulate the second set of constraints (3c). Again, due to Observation 3, for each $v,u\in T$, the inner maximization is
@@
$P_{\{v,u\}}(\mathcal{U}_{\text{SB}})=\mathcal{U}_{u,v}\equiv\{(\phi_{v},\phi_{u}):\quad\underaccent{\bar}{\phi}_{v}\leq\phi_{v}\leq\bar{\phi}_{v},\;\quad\underaccent{\bar}{\phi}_{u}\leq\phi_{u}\leq\bar{\phi}_{u},\;\quad\left|\phi_{v}-\phi_{u}\right|\leq\gamma_{vu}\}.$
@@
Let $u,v\in T$, and for convenience, define also $\tilde{\mathcal{U}}_{u,v}=\{(\phi_{v},\phi_{u}):-\gamma_{vu}\leq\phi_{v}-\phi_{u}\leq\gamma_{uv}\}$, to show that $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})\subseteq\mathcal{U}_{u,v}$. The contractive property of the projection onto a plane, as a convex set, implies that $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})\subseteq P_{u}(\mathcal{U}_{\text{SB}})\times P_{v}(\mathcal{U}_{\text{SB}})$. Then, by Proposition 4-(i) and the definition of $\mathcal{U}_{\text{SB}}$, it follows that
@@
Following the results of Propositions  4-5, formulation (3) with $\mathcal{U}=\mathcal{U}_{\text{SB}}$ can now be compactly reformulated with a finite number of constraints. In particular, in this reformulation, no variables are added and the number of constraints of this formulation is $2|T|$ times that of the nominal problem (1), thereby remaining polynomial in the size of the nominal formulation. Consider the formulation
$\underset{\begin{subarray}{c}{\bf x}\in\mathbb{R}^{n}_{+},\underaccent{\bar}{d}\in\mathbb{R}\end{subarray}}{\text{maximum}}\hskip-10.0pt$ $\underaccent{\bar}{d}$ (10a)
subject to $\underaccent{\bar}{\phi}_{v}d_{v}(x)\geq\underaccent{\bar}{d},$ $v\in T$ (10b)
$\bar{\phi}_{v}d_{v}(x)-\mu\max\{\bar{\phi}_{v}-\gamma_{vu},\underaccent{\bar}{\phi}_{u}\}d_{u}(x)\leq 0,\hskip-10.0pt$ $(u,v)\in T^{2}:u\neq v$ (10c)
$\min\{\underaccent{\bar}{\phi}_{u}+\gamma_{vu},\bar{\phi}_{v}\}d_{v}(x)-\mu\underaccent{\bar}{\phi}_{u}d_{u}(x)\leq 0,$ $(u,v)\in T^{2}:u\neq v$ (10d)
$\ d_{v}(x)\leq\bar{d}_{k},$ $k\in[K],v\in H_{k}.$ (10e)
The following theorem establishes the correctness of this formulation, and specifically, demonstrates that its optimal solution set coincides with that of (3).
@@
Suppose that Assumption 1 holds and let $\mathcal{U}=\mathcal{U}_{\text{SB}}$. Then, the optimal solution sets of (3) and (10) coincide.
@@
The reformulation of constraints (3b) is a straightforward consequence of (5) together and Proposition 4. Constraints (3c) are reformulated based on (8) to a maximization over $P_{\{v,u\}}(\mathcal{U}_{\text{SB}})$ in place of a maximization over $\mathcal{U}_{\text{SB}}$. Due to Assumption 1, this set is given by the result of Proposition 5. Further, since the objective is linear, and since, for each $u,v\in T$, $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})$ is a bound polyhedral set, then by the fundamental theorem of linear programming, there is an optimal $(\phi_{u},\phi_{v})$ that is an extreme point of the set $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})$. Moreover, since $\phi_{v}$ and $\phi_{u}$ have, respectively, a positive and a negative coefficient in (3c), this implies that the points in $\text{ext}(P_{\{u,v\}}(\mathcal{U}_{\text{SB}}))$ are dominated by the points in
@@
that is, for each pair $v,u\in T$, constraint (3c) is satisfied if and only if its left-hand side is nonpositive for these two points in $\mathcal{U}^{*}_{u,v}$. ∎
In Section 3, we will address the computational aspect of solving (10), specifically, how to deal with the potentially very large number of homogeneity constraints (10c) and (10d).
@@ 3 Solving Large-scale RO
The radiation therapy problems that we consider involve tens of thousands of tumor voxels and millions of healthy tissue voxels. Accordingly, formulation (10) may have hundreds of millions of constraints. Therefore, it may prove difficult to solve or even store this problem on most computers. In order to address this challenge, we propose an algorithm that dynamically generates constraints of (10) as needed.
@@
1 Phase I: Alternate between iteratively generating at most $n_{H}$ constraints  (10e) to each OAR until a feasibility tolerance $\tau_{\operatorname{Infs},1}$ is achieved and the objective function value does not decrease by more than a factor of $\tau_{z}$, and appending at most $n_{S}$ homogeneity constraints (10c) (10d) to set $S\subseteq T^{2}$. If none of the latter constraints are generated switch to Phase II;
Phase II: The same as Phase I, with feasibility tolerance of $\tau_{\operatorname{Infs},2}<\tau_{\operatorname{Infs},1}$ for constraints  (10e). Exit once no additional constraints can be generated.
@@
A detailed version of the algorithm in pseudocode is given as Algorithm 3 in Appendix A.1. The algorithm is initialized with beamlet configuration $x_{0}$, typically a part of a previously computed solution or an all-ones vector multiplied by some normalized dose. Based on this initial configuration, each OAR is initialized with the most violated $n_{0}$ constraints. This initialization, and the initial emphasis on generating OAR constraints (10e) in the algorithm, ensures that the optimization problem is bounded from the start or that it becomes bounded after a few iterations in phase I of the algorithm. The algorithm consists of two phases. Each one of them alternates between generation of OAR constraints (10e) and (10c) (10d). The first phase involves an OAR constraint infeasibility tolerance $\tau_{\operatorname{Infs},1}$ that is larger than the tolerance applied in the second phase, $\tau_{\operatorname{Infs},2}$. Intuitively, the larger tolerance in the first phase it intended to ensure both that the subproblems become bounded fairly quickly and to delay generation of OAR constraints that may be inactive at optimality until the homogeneity constraints are nearly satisfied.
@@ 4 Dose-Volume Constraints
In general, dose-volume constraints can be formulated by introducing additional auxiliary binary variables into formulation (10), turning it into the mixed-integer program,
@@
Lemma 7, together with the fact that there exists a value of $\beta>0$ for which $\sum_{v\in H_{K}}y_{v}(\beta)=0$, implies that for every value of $\Theta\geq 0$, it is possible to find a minimal value of $\beta$ such that (12c) is satisfied. Motivating our method is both monotonicity of $\left|\left|y(\beta)\right|\right|_{1}=\sum_{v\in H_{K}}y(\beta)_{v}$, following Lemma 7, and that the one-norm is a convex approximation of the zero-“norm” ($\left|\left|y(\beta)\right|\right|_{0}$) , which is the number of nonzero components of the deviation variable vector $y$. The one-norm is effectively used to promote sparsity in large-scale machine learning and compressed sensing applications (see, for example, Donoho and Elad, (2003)). We expect that solving (13) with appropriate $\beta$ will approximate the problem with dose-volume constraints (11f) reasonably well. Our goal is to find the minimal value of $\beta$ for which the optimal solution of (13) satisfies (11f). To achieve this , LP (13) is solved parametrically, starting from a lower bound on $\beta$ and traversing optimal bases for the given subsets of constraints defining the reduced problem. This algorithm is presented as Algorithm 2.
@@ 5 Experiments
In our experiments, we focus on a particular case study based on real brain data of Patient 4 from the TCIA archive (Clark et al., 2013), referred to as the Brain dataset in Table 1. Due to the very large number of constraints in (1), and especially in (10), for this dataset, we also use a smaller liver imaging dataset (Craft et al., 2014) to evaluate the computational performance of our dynamic generation of OAR and homogeneity constraints in Algorithm 3. FMISO-PET images are not available for this smaller dataset, so for this particular dataset, the radiosensitivity data are created synthetically. Oxygen levels tend to deccrease, and accordingly radiosensitivity values tend to decrease, as we approach the center of the PTV. Thus, to create sensible radiosensitivity values, we first define vector $f$ such that, for each PTV voxel $v\in T$, $f_{v}=(2\pi)^{-3/2}\det(\Sigma)^{-1/2}e^{-\frac{1}{2}(a_{v}-\overline{a})^{\top}\Sigma^{-1}(a_{v}-\overline{a})}$, where $a_{v}$ is the vector of x-y-z coordinates of voxel $v$’s position, $\overline{a}=\frac{1}{|T|}\sum_{v\in T}a_{v}$ is the mean voxel position, and the covariance is given by a ($3\times 3$) scaled covariance matrix of the PTV voxel positions, $\Sigma=\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}50}}{|T|}\sum_{v\in T}(a_{v}-\overline{a})(a_{v}-\overline{a})^{\top}$. The radiosensitivity for voxel $v\in T$ is then set to be $\phi_{v}=0.85+0.15\times\frac{f_{\max}-f_{v}}{f_{\max}-f_{\min}}$, where $f_{\max}=\max_{u\in T}\{f_{u}\}$ and $f_{\min}=\min_{u\in T}\{f_{u}\}$. This choice of $\phi_{v}$ implies that it is contained in $[0.85,1]$, while as voxel v’s position is closer to the center of the PTV, the radiosensitivity, $\phi_{v}$, approaches $0.85$. Details of both datasets are given in Table 1.
@@
$\left|T\right|$ $k$ Name $\left|H_{k}\right|$ $\bar{d}_{k}$ bounds [^1] Based on (ICR, 2010)
@@ 5.1 Computational Performance
Our implementation uses the Julia programming language, version 1.6.2, the JuMP package (Lubin and Dunning, 2015), and the Gurobi solver, version 9.1.2. The experiments are run on 40 core Linux servers each equipped with Intel Xeon E5-2650, 25 MB cache, 2.3 GHz base frequency CPUs, and 128 GB of RAM. Algorithm 1 (appearing in greater detail as Algorithm 3 in Appendix A.1) is invoked for solving the robust model (10) and the nominal model (1), with parameter values $\tau_{\operatorname{Infs},1}=10$, $\tau_{\operatorname{Infs},2}=0$, $\tau_{z}=10^{-2}$, and $n_{H}=400$. The homogeneity parameter is set to $\mu=1.15$ for Liver instances and $\mu=1.1250$ for Brain instances for the nominal experiments, and $\mu=1.1875$, which obtains a similar homogeneity and biologically-adjusted dose in the nominal scenario. Table 2 shows the running times achieved when solving the nominal model (1), which involves only $\left|T\right|$ homogeneity constraints. The results show that dynamic constraint generation is not needed to solve the nominal model but with careful tuning of the constraint generation algorithm, it can improve on the running time required to solve the full model (also yielding significant savings in terms of memory consumption), especially for larger data instances such as Brain. Table 3 shows the running times when applying the robust model to both datasets. The parameters of the uncertainty set are set to $\delta=0.04$ and $\gamma=0.04$ for Liver, and $\delta=0.08$ and $\gamma=0.04$ for Brain. Here, it is evident that dynamic constraint generation is necessary in order to be able to solve both models when the number of homogeneity constraints is $O(\left|T\right|^{2})$. However, dynamically generating all violated constraints may result in excessive running times. In the example shown in Table 3, dynamically generating all violated constraints in each iteration results in a running time of nearly four hours for the Brain instance. In preliminary experiments, with a different choice of parameters than experimented with here, when generating all violated constraints, we have observed cases where the running times took more than 48 hours to solve, and cases memory consumption exceeded its limits.
Table 2: Running time in seconds and number of constraints generated in solving the nominal model (1). Here, constraint generation, if indicated, implies that all violated constraints are generated in each iteration.
@@
Table 3: Running time in seconds and number of constraints generated in solving (10) using Algorithm 3.